The annual interest on a $19,000 investment exceeds the interest earned on a $12,000 investment by $531. The $19,000 is invested at a 0.4% higher rate of interest than the $12,000. What is the interest rate of each investment?
Answer by Boreal(15235)

interest rate on 12000 is x, so the interest is 12000x
for 19000, the interest rate is x+0.004, so the interest is 19000x+76
12000x+531=19000x+76
455=7000x
x=455/7000=0.065
interest is 12000*0.065=$780
x+0.004 is 0.069 or 6.9% interest rate and 19000*0.069=$1311
That is $531 higher.
rates are 6.5% and 6.9%